OpenStudy (anonymous):

x+ 2 / 15 + 3 - x / 5 = 2/3 x = 1 x = 0.5 x = 2

OpenStudy (mertsj):

\[\frac{x+2}{15}+\frac{3-x}{5}=\frac{2}{3}\]

OpenStudy (mertsj):

Multiply all terms by the common denominator which is 15

OpenStudy (anonymous):

I dont gt what u just said

OpenStudy (mertsj):

Which part do you not understand? Do you know what terms are?

OpenStudy (anonymous):

All of it, and no

OpenStudy (mertsj):